# Entanglement measures for detectability

### v2

@@ VII Algorithms
VII.1 Algorithm for $E^{*}(\rho_{AB})$
To calculate the proposed quantities for general states, we introduce algorithms for their calculation that can be applied to general states. Since the calculation of $E^{*}(\rho_{AB})$ is hard, we modify $E^{*}(\rho_{AB})$ as follows. We choose a finite set ${\cal D}$ of pure states on ${\cal H}_{B}$ as a discretization. That is, we denote ${\cal D}$ as $\{\rho_{B|c}^{*}\}_{c\in{\cal X}}$. Then, we define
@@ VII.1 Algorithm for $E^{*}(\rho_{AB})$
The above theorem is shown in Appendix C by using [18, Theorem 3.3].
VII.2 Algorithm for $E_{\alpha}^{*}(\rho_{AB})$
@@
4:until convergence.
@@
The above theorem is shown in Appendix C by using [18, Theorem 3.3].
@@ VIII Application of our algorithm
VIII.1 Membership problem
Next, we apply our algorithm to the membership problem of the set ${\cal S}_{AB}$. That is, we employ the quantity $E^{*}(\rho_{AB})$ for the membership problem. This measure will be used for constructing the separation hyperplane. In the following, we consider how to distinguish the following two cases.
@@ VIII.1 Membership problem
$t\geq\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}}}{\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))}$ (85)
@@
However, each step has various calculations. In each step, we need to calculate ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$. For this calculation, we need to calculate $\log\rho_{AC}$ and $\Omega_{R,D}[\rho_{AC}]$. $\rho_{AC}$ is a block diagonal matrix. That is, it is sufficient to calculate every component. Every component has a matrix function on a $d_{A}d_{B}$-dimensional space, and a factor whose complexity is linear to $|{\cal D}_{n,d_{B}}|$. But, the latter part is common for all $c$-components. The farmer can be done by diagonalization. The complexity of the diagonalization on $d$-dimensional matrix is $O(d^{3})$ with error $\epsilon$ [45, 46, 47]. Therefore, the calculation complexity of ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$ is $O(|{\cal D}_{n,d_{B}}|(d_{A}^{3}d_{B}^{3}))\leq O((\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))=O((\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))$. The total complexity is
@@
On the other hand, the method by [21] has the following calculation complexity
$O(d_{A}^{6}\epsilon^{-4d_{B}})$ (88)
@@
VIII.2 Separation hyperplane
Next, we discuss how to obtain separation hyperplane by using the above approximation and the geometrical characterization by Theorem 3. This is because it is a key step for conic linear programming with a general cone to find a separation hyperplane for an element that does not belong to the cone [25, 26, 27, 28]. For this aim, we employ the minimizer of $E_{{\cal D}}^{*}(\rho_{AB})$, i.e., the state
@@ IX Numerical analysis
For a typical example, we focus on the case when ${\mathcal{H}}_{A}$ is two qubits and ${\mathcal{H}}_{B}$ is one qubit. We parametrize a qubit pure state as $|\phi_{\theta}\rangle:=\frac{1}{\sqrt{2}}(|0\rangle+e^{\sqrt{-1}\theta}|1\rangle)$. Consider the entangled state $|\Psi_{p}\rangle:=\sqrt{p}|\phi_{0}\rangle|\phi_{0}\rangle|0\rangle+\sqrt{1-p}|\phi_{\pi/2}\rangle|\phi_{\pi/2}\rangle|1\rangle)$. Given a real number $\lambda$ with $10<\lambda<$ 1, we focus on the phase-damping channel ${\cal E}_{\lambda}$ defined as
@@
$\tilde{\rho}_{p,\lambda,\delta}:=(1-\delta)(\tilde{\cal E}_{\lambda}\otimes\tilde{\cal E}_{\lambda}\otimes Id)(|\tilde{\Psi}_{p}\rangle\langle\tilde{\Psi}_{p}|)+\delta\rho_{mix}.$ (109)
@@
Next, we apply our algorithm to the membership problem of separability by choosing $p=0.1$. Figure 4 plots the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=10^{-8}$. To compare our method with the partial transpose negative test, this figure plots the boundary to satisfy the condition that the minimum eigenvalue of the partial transposed state is $-10^{-8}$. These comparisons show that our method can detect entangled states that cannot be detected by the partial transpose negative test even when all the matrix components of the density matrix of a $4\times 2$ system are written in real numbers.
@@ X Exponential quantum Sanov theorem
We consider a $d$-dimensional Hilbert space ${\cal H}$ and a state $\rho$ on ${\mathcal{H}}$. We assume the $n$-fold independently and identical distributed (iid) condition and consider the following two hypotheses.
@@
$\lim_{n\to\infty}\frac{1}{n}-\log\beta_{\epsilon,n}(\rho)=E^{*}(\rho):=\min_{\sigma\in S}D(\sigma\|\rho).$ (113)
is known as quantum Sanov theorem [11, 12]. To get an exponential evaluation, as the extension of $E^{*}(\rho)$, we define the minimum sandwich relative entropy as
@@
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$ $\geq\sup_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}$ (115)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)$ $\leq\inf_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (116)
@@
Our bound is the optimal when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. Although the paper [12] derived exponential evaluation for both error probabilities, its exponent does not equal to the optimal bound even when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. In this sense, our evaluation is better than the upper bound by [12].
@@ XI Discussion
We have proposed new entanglement measures based on the distinguishability of the given entangled state from the set of separable states with two kinds of formulations by using quantum relative entropy and sandwich Rényi relative entropy. That is, we have proposed two kinds of measures based on quantum relative entropy. This measure is different from the entanglement of relative entropy [8, 9] because the choice of two input states are opposite to each other.
@@
Also, the proposed quantity is useful for finding entanglement witness based on its information-geometrical structure. In the next step, we have derived several useful calculation formulas for these measures. In particular, we have shown that these two kinds of measures coincide with each other when the entangled state is maximally correlated. In particular, the entanglement measure based on the quantum relative entropy has been concretely derived. However, it is not easy to calculate these measures for general entangled states. To resolve this problem, we have proposed algorithms to calculate these measures. In the first step of this calculation, we have approximated these values by introducing the $\delta$-net on the second quantum system, which was introduced to derive an entanglement witness [17, Section 3]. Then, we apply the quantum Arimoto-Blahut algorithm [18] to these approximating values. Then, we have obtained an efficient algorithm for these approximating values. These algorithms are shown to converge to the global minimum, and their errors are evaluated in a simple form.
@@ Acknowledgement
MH was supported in part by the National Natural Science Foundation of China under Grant 62171212. MH is thankful to Professor Hiroshi Hirai for helpful discussions on the membership problem for separability and the importance of separation hyperplane, and informing the references [15, 20, 25, 21, 23, 24]. MH is thankful to Professor Marco Tomamichel for informing the reference [22]. MH and YI are grateful for Professor Masaki Owari to a helpful discussion on membership problem of separability and informing the reference [17].
@@ Appendix A Construction of discrete subset
A.1 Real discrete subset
First, similar to [29, Section 5.4], we construct the real discrete subset ${\cal D}_{n,d,R}\subset{\cal S}_{d}$, where ${\cal S}_{d}$ is the $d$-dimensional sphare in $\mathbb{R}^{d+1}$. We define ${\cal D}_{n,1,R}$ as
@@ A.1 Real discrete subset
${\cal D}_{n,d,R}:=\Big{\{}\Big{(}\cos\frac{2\pi j}{n}v,\sin\frac{2\pi j}{n}\Big{)}\Big{|}v\in{\cal D}_{n,d-1,R},j=0,\ldots,n-1\Big{\}}.$ (119)
@@
This can be inductively shown as follows.
@@
A.2 Complex discrete subset
Next, we construct the complex dscrete subset ${\cal D}_{n,d}\subset\mathbb{C}^{d}$, where all elements of ${\cal D}_{n,d}$ are normalized vectors. Using ${\cal D}_{n,2d-2,R}$, we define ${\cal D}_{n,d}$ as
@@ Appendix B Proof of Lemma 7
Appendix C Proof of Theorems 8 and 9
C.1 Generalized Arimoto-Blahut algorithm
@@
The paper [18, Theorem 3.3] showed the following theorem, which discusses the convergence to the global minimum and the convergence speed.
@@ C.2 Proof of Theorem 8
$\Omega_{{\cal D}}[\rho_{AC}]:=\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (143)
@@
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\cal D}[\rho_{AC}]=\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{A}\sum_{c\in{\cal C}}X_{A|c}\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}=D\Big{(}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (144)
@@
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]=$ $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\exp(X_{A|c}-V_{c}[\{X_{A|c^{\prime}}\}_{c^{\prime}}]).$ (146)
@@
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}(\sum_{c\in{\cal C}}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}$
@@
$\leq$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\Big{)}=D(\rho_{AC}\|\rho_{AC}^{\prime}).$ (148)

### v3

@@ VII Algorithms
VII-A Algorithm for $E^{*}(\rho_{AB})$
Although the previous section presents the exact values of the proposed measures of maximally correlated states, it is not so easy to calculate the proposed measures of a general density matrix. To calculate the proposed quantities for general states, we introduce algorithms for their calculation that can be applied to general states. Since the calculation of $E^{*}(\rho_{AB})$ is hard, we modify $E^{*}(\rho_{AB})$ as follows. We choose a finite set ${\cal D}$ of pure states on ${\cal H}_{B}$ as a discretization. That is, we denote ${\cal D}$ as $\{\rho_{B|c}^{*}\}_{c\in{\cal X}}$. Then, we define
@@ VII-A Algorithm for $E^{*}(\rho_{AB})$
The above theorem is shown in Appendix C by using [28, Theorem 3.3].
VII-B Algorithm for $E_{\alpha}^{*}(\rho_{AB})$
@@
4:until convergence of $D_{\alpha}\Big{(}\sum_{c}X_{A|c}^{(t)}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}$.
@@
The above theorem is shown in Appendix C by using [28, Theorem 3.3].
@@ VIII Application of our algorithm
VIII-A Membership problem
The previous section presents algorithms to calculate the proposed measures. The special case of deciding if the entanglement measure is zero can be used for testing separability in a mathematical sense, which is called the membership problem of the set ${\cal S}_{AB}$. The aim of this section is the application of our algorithm to the membership problem of the set ${\cal S}_{AB}$. That is, we employ the quantity $E^{*}(\rho_{AB})$ for the membership problem. This measure will be used for constructing the separation hyperplane. In the following, we consider how to distinguish the following two cases.
@@ VIII-A Membership problem
$t\geq\frac{\log d_{A}+(d_{B}-1)\log\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}}}{\epsilon_{1}-(\epsilon_{2}(\|\log\rho_{AB}\|+\log d_{A}d_{B})+\eta_{0}(\epsilon_{2}))}.$ (93)
@@
However, each step has various calculations. In each step, we need to calculate ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$. For this calculation, we need to calculate $\log\rho_{AC}$ and $\Omega_{R,D}[\rho_{AC}]$. $\rho_{AC}$ is a block diagonal matrix. That is, it is sufficient to calculate every component. Every component has a matrix function on a $d_{A}d_{B}$-dimensional space, and a factor whose complexity is linear in $|{\cal D}_{n,d_{B}}|$. But, the latter part is common for all $c$-components. The former can be done by diagonalization. The complexity of the diagonalization on $d$-dimensional matrix is $O(d^{3})$ with error $\epsilon$ [64, 65, 66]. Therefore, the calculation complexity of ${\cal F}[\rho_{AC}]$ from $\rho_{AC}$ is $O(|{\cal D}_{n,d_{B}}|(d_{A}^{3}d_{B}^{3}))\leq O((\frac{2d_{B}^{1/2}\pi}{\epsilon_{2}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))=O((\frac{d_{B}^{1/2}\log d_{A}d_{B}}{\epsilon_{1}})^{d_{B}-1}(d_{A}^{3}d_{B}^{3}))$. The total complexity is
@@
On the other hand, the method by [31, Eq. (38)] has the following calculation complexity
$O(d_{A}^{6}\epsilon^{-4d_{B}}).$ (96)
@@
VIII-B Separation hyperplane, i.e., entanglement witness
A linear function to determine a separation hyperplane is called an entanglement witness in the context of entanglement theory when the cone is the separable cone. Next, we discuss how to obtain a separation hyperplane by using the above approximation and the geometrical characterization by Theorem 3. This is because it is a key step for conic linear programming with a general cone to find a separation hyperplane for an element that does not belong to the cone [35, 36, 37, 38]. For this aim, we employ the minimizer of $E_{{\cal D}}^{*}(\rho_{AB})$, i.e., the state
@@ IX Numerical analysis
Next, we numerically demonstrate our algorithm when ${\mathcal{H}}_{A}$ is two qubits and ${\mathcal{H}}_{B}$ is one qubit. We parametrize a qubit pure state as $|\phi_{\theta}\rangle:=\frac{1}{\sqrt{2}}(|0\rangle+e^{\sqrt{-1}\theta}|1\rangle)$. Consider the entangled state $|\Psi_{p}\rangle:=\sqrt{p}|\phi_{0}\rangle|\phi_{0}\rangle|0\rangle+\sqrt{1-p}|\phi_{\pi/2}\rangle|\phi_{\pi/2}\rangle|1\rangle)$. Given a real number $\lambda$ with $10<\lambda<$ 1, we focus on the phase-damping channel ${\cal E}_{\lambda}$ defined as
@@
$\tilde{\rho}_{p,\lambda,\delta}:=(1-\delta)\big{(}\tilde{\cal E}_{\lambda}\otimes\tilde{\cal E}_{\lambda}\otimes Id\big{)}(|\tilde{\Psi}_{p}\rangle\langle\tilde{\Psi}_{p}|)+\delta\rho_{mix}.$ (117)
@@
Next, we apply our algorithm to the membership problem of separability by choosing $p=0.1$. For this aim, ideally, it is needed to plot the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=0$. However, due to the limitation of the performance of our computer, we cannot handle a too small number. Hence, instead of the ideal value 00, we set the same small number $10^{-8}$. That is, Figure 4 plots the boundary to satisfy the condition $E_{{\cal D}}^{*}({\rho}_{0.1,\lambda,\delta})=10^{-8}$. To compare our method with the partial transpose negative test. This test also has the same problem to handle the case with zero eigenvalue. Instead of the ideal value 00, we set the same small number $-10^{-8}$ in the same way. That is, this figure plots the boundary to satisfy the condition that the minimum eigenvalue of the partial transposed state is $-10^{-8}$.
These comparisons show that our method can detect entangled states that cannot be detected by the partial transpose negative test even when all the matrix components of the density matrix of a $4\times 2$ system are written in real numbers.
@@ X Exponential quantum Sanov theorem
Finally, to clarify our contribution on quantum Sanov theorem, we state our obtained result for quantum Sanov theorem in this section. For this aim, we consider a $d$-dimensional Hilbert space ${\cal H}$ and a state $\rho$ on ${\mathcal{H}}$. We assume the $n$-fold independently and identical distributed (iid) condition and consider the following two hypotheses.
@@
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{\epsilon,n}(\rho)=E^{*}(\rho):=\min_{\sigma\in S}D(\sigma\|\rho)$ (121)
is known as quantum Sanov theorem [21, 22]. To analyze the case when the value $\epsilon$ goes to zero exponentially, as the extension of $E^{*}(\rho)$, we define the minimum sandwiched relative entropy as
@@
$\lim_{n\to\infty}-\frac{1}{n}\log\beta_{e^{-nr},n}(\rho)$ $\geq\max_{0\leq\alpha\leq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}$ (123)
$-\frac{1}{n}\log\beta_{1-e^{-nr},n}(\rho)$ $\leq\min_{\alpha\geq 1}\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}.$ (124)
@@
Our bound is the optimal when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. Although the paper [22] derived an exponential evaluation for both error probabilities, its exponent does not equal to the optimal bound even when the number of elements of $S$ is one and the unique element of $S$ is commutative with $\rho$. In this sense, our evaluation is better than the upper bound by [22].
@@ XI Discussion
We have proposed new entanglement measures based on the distinguishability of the given entangled state from the set of separable states with two kinds of formulations by using quantum relative entropy and sandwiched Rényi relative entropy. That is, we have proposed two kinds of measures based on quantum relative entropy. This measure is different from the relative entropy of entanglement [8, 9] because the choice of two input states are opposite to each other.
@@
Also, the proposed quantity is useful for finding entanglement witness based on its information-geometrical structure. In the next step, we have derived several useful calculation formulas for these measures. In particular, we have shown that these two kinds of measures coincide with each other when the entangled state is maximally correlated. In particular, the entanglement measure based on the quantum relative entropy has been concretely derived. However, it is not easy to calculate these measures for general entangled states. To resolve this problem, we have proposed algorithms to calculate these measures. In the first step of this calculation, we have approximated these values by introducing the $\delta$-net on the second quantum system, which was introduced to derive an entanglement witness [27, Section 3]. Then, we apply the quantum Arimoto-Blahut algorithm [28] to these approximating values. Then, we have obtained an efficient algorithm for these approximating values. These algorithms are shown to converge to the global minimum, and their errors are evaluated in a simple form.
@@ Acknowledgement
MH was supported in part by the National Natural Science Foundation of China under Grant 62171212. MH is thankful to Professor Hiroshi Hirai for helpful discussions on the membership problem for separability and the importance of separation hyperplane, and informing the references [25, 30, 35, 31, 33, 34]. MH is thankful to Professor Marco Tomamichel and Mr. Roberto Rubboli for informing the references [32], [55], [56]. MH is grateful to Dr. Baichu Yu for informing the references [46, 47]. MH and YI are grateful for Professor Masaki Owari to a helpful discussion on the membership problem of separability and informing the reference [27].
@@ Appendix A Construction of discrete subset
A-A Real discrete subset
First, similar to [39, Section 5.4], we construct the real discrete subset ${\cal D}_{n,d,R}\subset{\cal S}_{d}$, where ${\cal S}_{d}$ is the $d$-dimensional sphere in $\mathbb{R}^{d+1}$. We define ${\cal D}_{n,1,R}$ as
@@ A-A Real discrete subset
${\cal D}_{n,d,R}$
::=:= $\Big{\{}\Big{(}\cos\frac{2\pi j}{n}v,\sin\frac{2\pi j}{n}\Big{)}\Big{|}v\in{\cal D}_{n,d-1,R},j=0,\ldots,n-1\Big{\}}.$ (127)
@@
The relation (129) can be inductively shown as (130) in the next page.
@@
A-B Complex discrete subset
Next, we construct the complex discrete subset ${\cal D}_{n,d}\subset\mathbb{C}^{d}$, where all elements of ${\cal D}_{n,d}$ are normalized vectors. Using ${\cal D}_{n,2d-2,R}$, we define ${\cal D}_{n,d}$ as
@@ Appendix B Proof of Lemma 10
Appendix C Proof of Theorems 11 and 12
C-A Generalized Arimoto-Blahut algorithm
@@
The paper [28, Theorem 3.3] showed the following theorem, which discusses the convergence to the global minimum and the convergence speed.
@@ C-B Proof of Theorem 11
$\Omega_{{\cal D}}[\rho_{AC}]$
::=:= $\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (151)
@@
$\operatorname{\mathrm{Tr}}\rho_{AC}\Omega_{\cal D}[\rho_{AC}]$
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AC}(\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}})\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{A}\sum_{c\in{\cal C}}X_{A|c}\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}-\log\rho_{AB}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\log\rho_{AB}\Big{)}$
$=$ $D\Big{(}\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*}\Big{\|}\rho_{AB}\Big{)}.$ (152)
@@
${\cal F}[\sigma_{AC}[\{X_{A|c}\}_{c}]]$
$=$ $\frac{1}{\kappa[\{X_{A|c}\}_{c}]}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\exp(X_{A|c}-V_{c}[\{X_{A|c^{\prime}}\}_{c^{\prime}}]).$ (154)
@@
$=$ $\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\operatorname{\mathrm{Tr}}_{AB}(\sum_{c\in{\cal C}}P(c)\rho_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\log\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}$
$-\Big{(}\log\Big{(}\sum_{c^{\prime\prime}\in{\cal C}}X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}\Big{)}$
@@
$\leq$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes$
$X_{A|c^{\prime\prime}}^{\prime}\otimes\rho_{B|c^{\prime\prime}}^{*}\Big{)}$
$=$ $D\Big{(}\sum_{c^{\prime}\in{\cal C}}|c^{\prime}\rangle\langle c^{\prime}|\otimes X_{A|c^{\prime}}\Big{\|}\sum_{c^{\prime\prime}\in{\cal C}}|c^{\prime\prime}\rangle\langle c^{\prime\prime}|\otimes X_{A|c^{\prime\prime}}^{\prime}\Big{)}$
$=$ $D(\rho_{AC}\|\rho_{AC}^{\prime}).$ (156)